As a health physicist, you are being consulted aboPlease Help!
Answer #1

a)

dN/dt = 500 uCi = (500 x 10^-6)*(3.70 x 10^10 1/s) = 1.85 x 10^7 decay/sec.

T1/2 = ln(2) / \lambda

\lambda = ln(2) / T1/2

\lambda = ln(2) / 12d*(86400 s/d) = 6.69 x 10^-7 s

dN/dt = \lambda*N

N = (dN/dt)/\lambda

N = (6.69 x 10^7 sec) / (1.85 x 10^-7 decay/sec)

N = 2.77 x 10^13 nuclei

The mass of 131Ba nuclei is

m = 2.77 x 10^13 x (131 x 1.67 x 10^-27 kg/nucleus) = 6 x 10^-12 kg

m = 6 x 10^-9 gm

m = 6 ng.

b)

A = A0e^-\lambdat

1 uCi = (500 uCi)*e^-\lambdat

t = -ln(1/500) / \lambda

t = -ln(1/500) / (6.69 x 10^-7 1/sec)

t = 9.29 x 10^-6 sec (1d / 86400 sec)

t = 108 days